- Press Release
- Oct 4, 2022
NOAA GOES SATOPS Morning Report 17 November 2000
SATOPS MORNING REPORT November 17, 2000
No change in the status of the POES spacecraft.
At ~1100z to ~1447z yesterday morning, the FCDAS sustained an operational
power outage that affected FCDAS support activities during that period.
Multiple data losses were incurred (listed below), but significant manual
commanding and other “data chasing” activities by the crews minimized the
extent of the losses.
On Thursday, November 16 (rev 13060 / W at 2227z), the NOAA-15 HIRS
Filter Wheel Housing Heater was turned off as part of the continuing NOAA-15
HIRS Filter Wheel Test (Procedure # 001115.1525). Additional actions
today include the activation of the command MACRO to cycle the filter wheel
off for 2.5 sec durations during rev 13071 at 1719z and the turn-on of
the FW housing heater (precautionary for the weekend). Additional
activities next week include MACRO toggling at the FW low power mode.
NOAA-16 rev 791 / F at 1227z on November 16: L1A and G4A not recovered
due to FCDAS power outage. L1A recovery not possible due to recorder
management. G4A recovered on rev 794 / W.
NOAA-14 rev 30311 / F at 1341z on November 16: L5A and G2B not recovered
due to FCDAS power outage. L5A recovery not possible due to recorder
management. G1B recovered on rev 30312 / W.
NOAA-16 rev 792 / F at 1408z on November 16: L2A and G5B not recovered
due to FCDAS power outage. L2A recovery not possible due to recorder
management. G5B recovered on rev 794 / W.
NOAA-10 rev 73653 / F at 1426z on November 16: S3B late in locking on
to data (first FCDAS acquisition since the power outage). 31.0% of
NOAA-14 rev 30312 / F at 1521z on November 16: G1B recovered from
rev 30311. L2B not recovered due to time constraints. L2B recovery
not possible due to recorder management.
NOAA-16 rev 794 / W at 1712z on November 16: Recovered G5B (Rev
# 792F) and G4A (Rev # 793F).
NOAA-16 rev 796 / F at 2046z on November 16: Executed CP to change
Solar Array offset from -40 deg to -35 degrees.
POES operations were otherwise nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
At 2102z on Thursday, November 16, the GOES-10 IMC upload was loaded
late (due to an initial load failure). As a result, the 2120z sounding
GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at ~105 deg W.
GOES operations were nominal over the past 24 hours.
Questions or comments to firstname.lastname@example.org